OpenStudy (anonymous):

how do you find n in this type of equation: 81=3 to the power of n-1? Obviously here n is 5, but is there some formula for figuring that out?

OpenStudy (anonymous):

81 = 3^4 4 = n-1 n = 5

OpenStudy (anonymous):

yeah, i get that, but assuming that the #s aren't 81 and 3, and you don't already know what the exponent should be. How would you figure it out?

OpenStudy (anonymous):

You can use logarithum log 81 / log 3 = n-1

OpenStudy (anonymous):

thanks